# OncoReg: Medical Image Registration for Oncological Challenges

### v1

@@ Abstract
This work details the methodology and data behind the OncoReg Challenge and provides a comprehensive analysis of the competition entries and results. Findings reveal that feature extraction plays a pivotal role in this registration task. A new method emerging from this challenge demonstrated its versatility, while established approaches continue to perform comparably to newer techniques. Both deep learning and classical approaches still play significant roles in image registration, with the combination of methods—particularly in feature extraction—proving most effective.
Keywords: Machine Learning, Image Registration, Cone-beam Computed Tomography, Image Guided Radiotherapy
@@ 1 Introduction
In radiology, and particularly in modern cancer research, vast amounts of data are continuously generated. However, much of this data remains underutilized for developing innovative AI-centered solutions due to several challenges. One significant obstacle is the restriction on data publication imposed by patient privacy rights. This is especially problematic for data collected in previous years, where obtaining retroactive patient consent is no longer feasible. To unlock the potential of such data while ensuring compliance with privacy laws, novel approaches are required.
The OncoReg Challenge addresses this issue by enabling researchers to develop and validate methods in a two-phase framework. In the first phase, challenge participants work with a publicly available dataset to design and refine their algorithms. In the second phase, the submitted models are trained in-house on a private, unpublished dataset—ensuring that the data remains securely within the hospital network and never becomes accessible to external participants.
This approach not only safeguards patient privacy but also fosters the development of more generalizable AI models. Since participants must optimize their methods for unseen data rather than a specific dataset, the challenge encourages the creation of algorithms with broader applicability and better real-world transferability. Computational challenges provide an ideal framework for this kind of data utilization, allowing multiple research groups to benefit from the dataset without the need for public distribution. The popularity of challenges, especially in the field of computer vision, has been steadily increasing, as reflected in the growing number of challenges hosted at major conferences such as MICCAI, ISBI, and SPIE Medical Imaging.
In the field of medical image registration, the Learn2Reg Challenge (Hering et al., 2022) holds a prime position. The well-known challenge, which has been held annually since 2019 in conjunction with MICCAI, has significantly advanced algorithmic approaches and benchmarking in medical image registration, fostering collaboration and innovation among researchers worldwide. Learn2Reg offers an extensive selection of 3D image registration tasks, covering a variety of scenarios. The tasks include both mono- and multimodal registration as well as inter- and intrapatient registration. The challenges span multiple imaging modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and ultrasound (US), and encompass different anatomical regions, including the lungs, abdomen, and brain. OncoReg builds upon the foundation established by Learn2Reg, extending it to include an additional modality—cone-beam CT—and shifting the focus to real-world oncological challenges. The first part of OncoReg comprises a task (ThoraxCBCT) in the Learn2Reg Challenge 2023, followed by a standalone virtual challenge and workshop.
@@
In particular, the registration of low-dose interventional cone-beam CTs (CBCTs) with high-resolution diagnostic fan-beam CTs (FBCTs) is a fundamental step in image-guided radiotherapy. FBCTs, acquired before treatment, provide a detailed map of the patient’s anatomy and are used to plan the radiation dose distribution. However, during the course of treatment, the patient’s anatomy can change due to tumour shrinkage, weight loss, or other factors. CBCTs are frequently acquired during treatment sessions to assess these changes in real time. By registering CBCTs with the initial planning CTs, clinicians can adjust the treatment plan dynamically, ensuring that radiation continues to be delivered precisely to the target area, even as the anatomy evolves.
@@ 2.2 Cone-Beam CT Registration
Despite the widespread use of CBCT, its registration to conventional CT remains a challenging problem due to lower image quality, artifacts, and intensity variations. While general-purpose registration methods have been evaluated on CBCT-CT tasks (Demir et al., 2024; van Harten et al., 2024), dedicated solutions remain relatively scarce. Historically, semi-automatic optimization-based methods have been applied (Sotiras et al., 2013), and classical optimization approaches continue to be refined (Xie et al., 2022). More recently, deep learning-based methods have emerged, such as an unsupervised dual-attention network (Hu et al., 2022) and a rigid deep-learning-based registration method for head-neck CT-CBCT (Peng et al., 2024). Additionally, (Teng et al., 2021) introduced a supervised CNN for phase-to-phase registration in 4D CBCT, and (Li et al., 2023) proposed generating high-quality synthetic CTs from CBCTs as an alternative approach to facilitate registration.
@@ 3.1 Challenge Organisation
Computational challenges in the field of machine learning are often categorised based on the accessibility of data to participants. These challenges generally fall into four distinct types. Type 0 challenges offer participants access to all data phases, including supervision data, which minimises the workload for organisers. However, this approach compromises fairness, as participants are responsible for computing their own metrics, leading to potential inconsistencies. Type 1 challenges improve on this by keeping the supervision data for the test phase private, while the rest of the data is publicly accessible. Participants submit their predictions, which the organisers then evaluate and rank. This structure ensures a more consistent evaluation process but still allows participants access to much of the data. Type 2 challenges further increase the complexity by withholding not just the supervision data but also the test phase data itself. Participants must submit an inference algorithm, which the organisers run on the hidden data before evaluating and ranking the results. This approach increases the workload for both organisers and participants. Participants often need to containerise their methods or compile them into specific formats, adding to the complexity. Type 3 challenges represent the most complex scenario, where both the training and test data are hidden. Participants must submit their entire training algorithm, which organisers then use to retrain models on the hidden data before conducting inference, evaluation, and ranking. While this increases the burden on organisers significantly, it offers several advantages. One key benefit is that sensitive data—such as medical data that cannot be publicly disclosed due to patient privacy concerns—can still be utilised for research and algorithm development. Additionally, this approach ensures that all submissions are evaluated under identical conditions, eliminating the use of additional data, annotations, or varied hardware, thereby drastically enhancing fairness.
@@
3.1.1 Learn2Reg 2023 ThoraxCBCT Task
In 2023, the Learn2Reg challenge expanded its scope by introducing the new ThoraxCBCT task and extending the National Lung Screening Trial (NLST) task from the 2022. It also continued previous tasks such as LungCT, OASIS, and AbdomenMRCT, building on the challenge’s comprehensive history (Hering et al., 2022). The two primary tasks of the competition were hosted entirely on grand-challenge.org and followed a structured, two-phase process:
@@ 3.1.2 OncoReg Workshop
3.1.3 Lowering Entry Barriers
@@
3.2 Tasks & Data
3.2.1 ThoraxCBCT
@@
The dataset released includes training images from 11 patients, comprising 11 FBCT and 22 CBCT scans, and validation images from 3 patients, which consist of 3 FBCT and 6 CBCT scans. Each patient study includes two image pairs, resulting in a total of 22 training pairs and 6 validation pairs. For every image pair, Förstner keypoints are provided, based on displacement fields computed using ConvexAdam (Siebert et al., 2021). Unlike previous Learn2Reg tasks that focused primarily on the lungs, this task extends to include other thoracic structures, particularly organs at risk in radiotherapy. Keypoints were therefore determined not only within the lung but across the entire trunk region.
The test dataset includes images from 5 patients, yielding 5 FBCT and 10 CBCT scans (10 test pairs). The validation and test annotations feature multiple segmentation masks for various structures, including lungs, tumors, organs at risk (heart, spinal cord, esophagus, trachea, and aorta), and skeletal structures such as vertebrae, ribs, clavicles, and scapulae. Additionally, anatomical landmarks like airway bifurcations, the top of the aortic arch, apex cordis, and key points in bone structures were marked. Some annotations for FBCTs were sourced from TCIA, with missing segmentation masks generated using the TotalSegmentator tool (Wasserthal et al., 2023) and propagated to CBCTs through registration. These generated masks were validated by clinicians at the University Hospital of Schleswig-Holstein (UKSH). On average, 13 structures were segmented per image. Landmark annotations were entirely manual, resulting in an average of 16 landmarks per image pair. A binary body-trunk mask, distinguishing the thorax from the background, was provided for each image, also created using the TotalSegmentator.
@@
3.2.2 OncoReg
The OncoReg task extends the ThoraxCBCT challenge but is conducted as a type 3 task using a completely hidden dataset. Participants are tasked with registering a high-resolution diagnostic FBCT acquired before therapy with two CBCT scans: one from the start and another from the end of therapy. Unlike the ThoraxCBCT data, OncoReg images are not synchronized with respect to breathing phases, meaning the scans may capture different states of inspiration and expiration. The data, collected from routine clinical care at the University Hospital of Schleswig-Holstein (UKSH), was retrospectively curated from 25 lung cancer patients and carefully annotated with semi-automatic segmentations and manual landmarks. Similar to ThoraxCBCT, trunk masks and Förstner keypoints were computed for each pair. Preprocessing included converting images from DICOM to NIfTI format, cropping, resampling to a size of 256x192x192, with a voxel spacing of 1.5x1.5x1.5 mm, and manual pre-alignment of image pairs. Annotations for OncoReg were more comprehensive than those for ThoraxCBCT, as the entire dataset was annotated, rather than just a subset. Segmentations were created using the TotalSegmentator tool and then manually refined, resulting in an average of 52 annotated structures per image. Manual landmark annotation was also more detailed, with an average of 25 landmarks identified per image pair. While this dataset is not publicly available, researchers may request evaluation of their registration methods on the data via the challenge platform.
@@ 3.3 Challenge Design
To ensure a comprehensive evaluation of the registration performance, a range of complementary metrics are employed that are adapted after previous editions of Learn2Reg and are outlined in Table 1. These metrics are chosen to evaluate various aspects of algorithmic performance, including accuracy, robustness, plausibility, and computational efficiency. The final ranking also accounts for the statistical significance of the differences between results, using a ranking system inspired by the Medical Decathlon (Antonelli et al., 2022). Methods are statistically compared using the Wilcoxon signed-rank test, with a significance level set at $p<0.05$. Based on the number of “won” comparisons, each method receives a numerical rank score ranging from 0.1 to 1. The overall task ranking is then determined by calculating the geometric mean of these individual metric scores.
@@
3.4 Challenge Entries
The ThoraxCBCT validation leaderboard featured 68 entries from seven teams (including the challenge organisers), with seven valid submissions made during the test phase. Four of these submissions were baseline methods provided by the organisers.
The type 3 OncoReg task received five submissions from four participating teams, in addition to five baseline methods supplied by the organisers. The baseline methods included Niftyreg (Modat et al., 2009), DeedsBCV (Heinrich et al., 2013, 2015), ConvexAdam (Siebert et al., 2022; Heinrich et al., 2014), and Voxelmorph++ (Heinrich and Hansen, 2022). Voxelmorph++, in particular, was made available as an algorithm on grand-challenge.org/algorithms and was containerised as an example on the challenge’s GitHub repository, which is why it is discussed in detail in the following section. Additionally, a keypoint based registration method using the medical open network for AI (MONAI) (Cardoso et al., 2022) is provided. The baseline methods will not be further elaborated upon here, as they are already well-documented in existing literature. The methods submitted by the participating teams are described in detail below.
3.4.1 DINO-Reg
@@ 3.4.2 Fourier-Net
Image Augmentation: Online data augmentation was applied during pre-training on the ThoraxCBCT dataset. Specifically, images were randomly rotated, cropped, and subjected to noise addition, using the default probability settings provided by https://github.com/ZFTurbo/volumentations. The training process on the private dataset was similar to that used for the ThoraxCBCT dataset, where image similarity, keypoint mask Dice, and bending energy were optimised. Online data augmentation was also employed during this stage. The instance-level optimisation involved fine-tuning the trained weights on each test pair. For this challenge, the number of iterations was fixed at 700, with a learning rate of 0.0001 for each test pair.
3.4.3 SynDeeds
The OncoReg task focuses on the challenging yet clinically crucial objective of aligning pre-therapeutic, high-quality fan-beam CT (FBCT) with interventional low-dose cone-beam CT (CBCT) for image-guided radiation therapy. The low contrast resolution and frequent presence of metal artifacts in CBCT images create significant difficulties in achieving accurate registration.

### v3

@@ Abstract
This work details the methodology and data behind the OncoReg Challenge and provides a comprehensive analysis of the competition entries and results. Findings reveal that feature extraction plays a pivotal role in this registration task. A new method emerging from this challenge demonstrated its versatility, while established approaches continue to perform comparably to newer techniques. Both, deep learning and classical approaches still play significant roles in image registration, with the combination of methods, particularly in feature extraction, proving most effective.
doi:
10.59275/j.melba.2024-AAAA
keywords:
Image Registration, Cone-beam Computed Tomography, Image Guided Radiotherapy, Computational Challenge
volume: 2
@@ 1 Introduction
[…]
In radiology, and particularly in modern cancer research, vast amounts of data are continuously generated. However, much of this data remains underutilised for developing innovative AI-centered solutions due to several challenges. One significant obstacle is the restriction on data publication imposed by patient privacy rights. This is especially problematic for data collected in previous years, where obtaining retroactive patient consent is no longer feasible. To unlock the potential of such data while ensuring compliance with privacy laws, novel approaches are required. The OncoReg Challenge addresses this issue by enabling researchers to develop and validate methods in a two-phase framework. In the first phase, challenge participants work with a publicly available dataset to design and refine their algorithms. In the second phase, the submitted models are trained in-house on a private, unpublished dataset—ensuring that the data remains securely within the hospital network and never becomes accessible to external participants.
This approach not only safeguards patient privacy but also fosters the development of more generalisable AI models. Since participants must optimise their methods for unseen data rather than a specific dataset, the challenge encourages the creation of algorithms with broader applicability and better real-world transferability. Computational challenges provide an ideal framework for this kind of data utilisation, allowing multiple research groups to benefit from the dataset without the need for public distribution. The popularity of challenges, especially in the field of computer vision, has been steadily increasing, as reflected in the growing number of challenges hosted at major conferences such as MICCAI, ISBI, and SPIE Medical Imaging.
In the field of medical image registration, the Learn2Reg Challenge (Hering et al., 2022) holds a prime position. The well-known challenge, which has been held annually since 2019 in conjunction with MICCAI, has significantly advanced algorithmic approaches and benchmarking in medical image registration, fostering collaboration and innovation among researchers worldwide. Learn2Reg offers an extensive selection of 3D image registration tasks, covering a variety of scenarios. The tasks include both mono- and multimodal registration as well as inter- and intrapatient registration. The challenges span multiple imaging modalities such as computed tomography (CT), magnetic resonance imaging (MRI), and ultrasound (US), and encompass different anatomical regions, including the lungs, abdomen, and brain. OncoReg builds upon the foundation established by Learn2Reg, extending it to include an additional modality, cone-beam CT, and shifting the focus to real-world oncological challenges. The first part of OncoReg comprises a task (ThoraxCBCT) in the Learn2Reg Challenge 2023, followed by a standalone virtual challenge and workshop.
@@
In particular, the registration of low-dose interventional cone-beam CTs (CBCTs) with high-resolution diagnostic fan-beam CTs (FBCTs) is a fundamental step in image-guided radiotherapy (Zhang et al., 2024; Cao et al., 2022; Han et al., 2021). FBCTs, acquired before treatment, provide a detailed map of the patient’s anatomy and are used to plan the radiation dose distribution. However, during the course of treatment, the patient’s anatomy can change due to tumour shrinkage, weight loss, or other factors. CBCTs are frequently acquired during treatment sessions to assess these changes in real time. By registering CBCTs with the initial planning CTs, clinicians can adjust the treatment plan dynamically, ensuring that radiation continues to be delivered precisely to the target area, even as the anatomy evolves.
@@ 2.2 Cone-Beam CT Registration
Despite the widespread use of CBCT, its registration to conventional CT remains a challenging problem due to lower image quality, artifacts, and intensity variations. While general-purpose registration methods have been evaluated on CBCT-CT tasks (Demir et al., 2024; van Harten et al., 2024), dedicated solutions remain relatively scarce. Historically, semi-automatic optimisation-based methods have been applied (Sotiras et al., 2013), and classical optimisation approaches continue to be refined (Xie et al., 2022). More recently, deep learning-based methods have emerged, such as an unsupervised dual-attention network (Hu et al., 2022) and a rigid deep-learning-based registration method for head-neck CT-CBCT (Peng et al., 2024). Additionally, (Teng et al., 2021) introduced a supervised CNN for phase-to-phase registration in 4D CBCT, and (Li et al., 2023) proposed generating high-quality synthetic CTs from CBCTs as an alternative approach to facilitate registration.
@@ 3 Challenge Organisation
Computational challenges in the field of machine learning are often categorised based on the accessibility of data to participants. These challenges generally fall into four distinct types. Type 0 challenges offer participants access to all data phases, including supervision data, which minimises the workload for organisers. However, this approach compromises fairness, as participants are responsible for computing their own metrics, leading to potential inconsistencies. Type 1 challenges improve on this by keeping the supervision data for the test phase private, while the rest of the data is publicly accessible. Participants submit their predictions, which the organisers then evaluate and rank. This structure ensures a more consistent evaluation process but still allows participants access to much of the data. Type 2 challenges further increase the complexity by withholding not just the supervision data but also the test phase data itself. Participants must submit an inference algorithm, which the organisers run on the hidden data before evaluating and ranking the results. This approach increases the workload for both organisers and participants. Participants often need to containerise their methods or compile them into specific formats, adding to the complexity. Type 3 challenges represent the most complex scenario, where both the training and test data are hidden. Participants must submit their entire training algorithm, which organisers then use to retrain models on the hidden data before conducting inference, evaluation, and ranking. While this increases the burden on organisers significantly, it offers several advantages. One key benefit is that sensitive data, such as medical data that cannot be publicly disclosed due to patient privacy concerns, can still be utilised for research and algorithm development. Additionally, this approach ensures that all submissions are evaluated under identical conditions, eliminating the use of additional data, annotations, or varied hardware, thereby drastically enhancing fairness.
@@
3.1 Learn2Reg 2023 ThoraxCBCT Task
In 2023, the Learn2Reg challenge expanded its scope by introducing the new ThoraxCBCT task and extending the National Lung Screening Trial (NLST) task from the 2022 iteration. It also continued previous tasks such as LungCT, OASIS, and AbdomenMRCT, building on the challenge’s comprehensive history (Hering et al., 2022). The two primary tasks of the competition were hosted entirely on grand-challenge.org and followed a structured, two-phase process:
@@ 3.2 OncoReg Workshop
3.3 Lowering Entry Barriers
@@
4 Tasks & Data
Table 1: Overview of the two datasets used for the ThoraxCBCT and OncoReg tasks.
Dataset ThoraxCBCT OncoReg
Modality Fan-beam CT & Cone-beam CT Fan-beam CT & Cone-beam CT
4D (max. Insp./Exp.) $\times$
# Cases (2 Pairs/Case) 19 (1xFBCT, 2xCBCT per Case) 27 (1xFBCT, 2xCBCT per Case)
Split 11 Train, 3 Val., 5 Test 3-Fold CV (15 Train, 3 Val., 9 Test)
Image Size 390, 280, 300 256, 192, 192
Spacing 1 mm isotrope 1.5 mm isotrope
Manual Pre-alignment
Label Masks Semi-automatic, 13 Structures Semi-automatic, 52 Structures
Landmark Annotations Manual, Avg. 16 per Image Manual, Avg. 25 per Image
Binary Trunc Masks
Paired Förstner Keypoints ✓(DeedsBCV) ✓(ConvexAdam)
Access Original Data publicly available via TCIA, Processed Data available on Challenge Website Private Dataset, Evaluation on data per request
4.1 ThoraxCBCT
@@
The dataset released includes training images from 11 patients, comprising 11 FBCT and 22 CBCT scans, and validation images from 3 patients, which consist of 3 FBCT and 6 CBCT scans. Each patient study includes two image pairs, resulting in a total of 22 training pairs and 6 validation pairs. For every image pair, Förstner keypoints are provided, based on displacement fields computed using DeedsBCV (Heinrich et al., 2013). Unlike previous Learn2Reg tasks that focused primarily on the lungs, this task extends to include other thoracic structures, particularly organs at risk in radiotherapy. Keypoints were therefore determined not only within the lung but across the entire trunc region.
The test dataset includes images from 5 patients, yielding 5 FBCT and 10 CBCT scans (10 test pairs). The validation and test annotations feature multiple segmentation masks for various structures, including lungs, tumors, organs at risk (heart, spinal cord, esophagus, trachea, and aorta), and skeletal structures such as vertebrae, ribs, clavicles, and scapulae. Additionally, anatomical landmarks like airway bifurcations, the top of the aortic arch, apex cordis, and key points in bone structures were marked. Some annotations for FBCTs were sourced from TCIA, with missing segmentation masks generated using the TotalSegmentator tool (Wasserthal et al., 2023) and propagated to CBCTs through registration. These generated masks were validated by clinicians at the University Hospital of Schleswig-Holstein (UKSH). On average, 13 structures were segmented per image. Landmark annotations were entirely manual, resulting in an average of 16 landmarks per image pair. A binary body-trunc mask, distinguishing the thorax from the background, was provided for each image, also created using the TotalSegmentator. An example for an image pair of the ThoraxCBCT dataset including label mask annotations is given in Fig. 1.
@@
Figure 1: Exemplary image pair from the ThoraxCBCT dataset including semi-automatic label mask annotations. Left: Fan-Beam CT. Right: Cone-Beam CT.
4.2 OncoReg
The OncoReg task extends the ThoraxCBCT challenge but is conducted as a type 3 task using a completely hidden dataset. Participants are tasked with registering a high-resolution diagnostic FBCT acquired before therapy with two CBCT scans: one from the start and another from the end of therapy. Unlike the ThoraxCBCT data, OncoReg images are not synchronized with respect to breathing phases, meaning the scans may capture different states of inspiration and expiration. The data, collected from routine clinical care at the University Hospital of Schleswig-Holstein (UKSH), was retrospectively curated from 25 lung cancer patients and carefully annotated with semi-automatic segmentations and manual landmarks. Similar to ThoraxCBCT, trunc masks and Förstner keypoints were computed for each pair. Preprocessing included converting images from DICOM to NIfTI format, cropping, resampling to a size of 256x192x192, with a voxel spacing of 1.5x1.5x1.5 mm, and manual pre-alignment of image pairs.
Annotations for OncoReg were more comprehensive than those for ThoraxCBCT, as the entire dataset was annotated, rather than just a subset. Segmentations were created using the TotalSegmentator tool and then manually refined, resulting in an average of 52 annotated structures per image. Manual landmark annotation was also more detailed, with an average of 25 landmarks identified per image pair. An overview of the label mask annotations is given in Fig. 2.
While this dataset is not publicly available, researchers may request evaluation of their registration methods on the data via the challenge platform. An overview of the two datasets is given in Table 1.
Figure 2: Semi-automatic label mask annotations for an exemplary Fan-Beam CT from the OncoReg dataset. 3D rendering generated with 3D Slicer (Fedorov et al., 2012).
@@ 4.3 Challenge Design
To ensure a comprehensive evaluation of the registration performance, a range of complementary metrics are employed that are adapted after previous editions of Learn2Reg and are outlined in Table 2. These metrics are chosen to evaluate various aspects of algorithmic performance, including accuracy, robustness, plausibility, and computational efficiency. The final ranking also accounts for the statistical significance of the differences between results, using a ranking system inspired by the Medical Decathlon (Antonelli et al., 2022). Methods are statistically compared using the Wilcoxon signed-rank test, with a significance level set at $p<0.05$. Based on the number of won comparisons, each method receives a numerical rank score ranging from 0.1 to 1. The overall task ranking is then determined by calculating the geometric mean of these individual metric scores.
@@
5 Challenge Entries
The ThoraxCBCT validation leaderboard featured 68 entries from seven teams (including the challenge organisers), with seven valid submissions made during the test phase. Four of these submissions were baseline methods provided by the organisers. The type 3 OncoReg task received five submissions from four participating teams, in addition to five baseline methods supplied by the organisers. The baseline methods included Niftyreg (Modat et al., 2009), DeedsBCV (Heinrich et al., 2013, 2015), ConvexAdam (Siebert et al., 2021; Heinrich et al., 2014), and Voxelmorph++ (Heinrich and Hansen, 2022). Voxelmorph++, in particular, was made available as an algorithm on grand-challenge.org/algorithms and was containerised as an example on the challenge’s GitHub repository, which is why it is discussed in detail in the following section. Additionally, a keypoint based registration method using the medical open network for AI (MONAI) (Cardoso et al., 2022) is provided. The baseline methods will not be further elaborated upon here, as they are already well-documented in existing literature. The methods submitted by the participating teams are described in detail below.
5.1 DINO-Reg & DINO-Reg Ensemble
@@ 5.2 Fourier-Net
Image Augmentation: Online data augmentation was applied during pre-training on the ThoraxCBCT dataset. Specifically, images were randomly rotated, cropped, and subjected to noise addition, using the default probability settings provided by https://github.com/ZFTurbo/volumentations.
The training process on the private dataset was similar to that used for the ThoraxCBCT dataset, where image similarity, keypoint mask Dice, and bending energy were optimised. Online data augmentation was also employed during this stage. The instance-level optimisation involved fine-tuning the trained weights on each test pair. For this challenge, the number of iterations was fixed at 700, with a learning rate of 0.0001 for each test pair.
5.3 SynDeeds